OpenStudy (anonymous):

A man stands on the roof of a building of height 13.0m and throws a rock with a velocity of magnitude 34.5m/s at an angle of 26.1^\circ above the horizontal. You can ignore air resistance. calculate the horizontal distance from the base of the building to the point where the rock hits the ground?

OpenStudy (anonymous):

\[x=vt \cos \theta \] we need to get tıme \[y=vt \sin \theta+\frac{ 1 }{ 2 }at ^{2}\] \[-13=(t)34,5\sin (21,6)+\frac{ 1 }{ 2 }(-9.8)t^2\] \[4.9t^2-t(34.5 \sin 21.6)-13=0\] use quadratıc

OpenStudy (anonymous):

\[t=3.377425129\] \[x=34.5 \cos 21.6 \times 3.37\] \[x=108.3386411\]